Could you please help me with the following problem. Thank you in advance.
An overzealous student burns the candle at both ends. One night he lights
the top and the bottom of a 10 inch candle at the same time. If the top
flame burns at a speed of 1.4 in/hr. and the bottom flame burns at a rate
of 2.1 in/hr., how far from the top do the flames meet?

Answer by ankor@dixie-net.com(22740)

the top and the bottom of a 10 inch candle at the same time. If the top
flame burns at a speed of 1.4 in/hr. and the bottom flame burns at a rate
of 2.1 in/hr., how far from the top do the flames meet?
:
Here's one way to do it:
:
Find how long it takes for this to occur
:
Let t = time required for the two flames to meet:
:
1.4t + 2.1t = 10
3.5t = 10
t = 10/3.5
t = 2.857 hrs
:
Dist from top:
2.847 * 1.4 = 4" from top
or
2.857 * 2.1 = 6" from the bottom